A balloon of total mass m is descending with an acceleration of a. Find the mass that removed from the balloon such that it ascends with the same acceleration a.
Answer must be in terms of a,g(accelearion due to gravity) and m

Grade:10

1 Answers

AAYUSH KESHAVA
11 years ago

given

mg-fair=ma                   (i)

acording to the question;

fair-(m-x)g=(m-x)a    (ii)

(i)+(ii) gives us
g(m-m+x)=a(m+m-x)

gx/a=2m-x

x(g/a+1)=2m

solving for x immidiately gives us

x(g+a)/a=2m

x=(2ma)/(g+a)